# Linear-scaling Tensor Train Sketching

### 11 Mar 2026 v1

@@ 6.2 Properties of Gaussian Tensor-Train Sketches
By Lemma 6.3, condition (13) ensures that the Gaussian matrices $\mathbf{G}_{j}$ satisfy the Strong $(\varepsilon/L\sqrt{d},\delta)$-JLM Property and by Imported Lemma 6.6, so do matrices $\mathbf{M}_{j}=\mathbf{I}_{n_{1}\cdots n_{j-1}}\otimes\mathbf{G}_{j}$. Finally, by Imported Lemma 6.7, the product ${\bm{\Omega}}=\mathbf{M}_{1}\cdots\mathbf{M}_{d}$ satisfies the Strong $(\varepsilon,\delta)$-JLM Property. ∎
@@
6.3 Properties of Block-Sparse Tensor Train sketches
Now knowing that each row block of the BSTT sketch is indeed a subspace embedding for large enough $R$, we can proceed to showing that this unifying framework is also an OSE, where the number of blocks $P$ helps relax the constraints on the ranks $R$. To this effect, we show the following result:
@@ 6.3 Properties of Block-Sparse Tensor Train sketches
There exists a universal constant $c>0$ such that for iid ${\bm{\Omega}}_{1},\dots,{\bm{\Omega}}_{P}$ iid realizations of a random matrix with the Strong $\bigl(\frac{\sqrt{P}\varepsilon}{4e^{2}},\delta\bigr)$-JLM property, then the random matrix ${\bm{\Omega}}=\frac{1}{\sqrt{P}}\begin{bmatrix}{\bm{\Omega}}_{1}\\ \vdots\\ {\bm{\Omega}}_{P}\end{bmatrix}$ has the Strong $(\varepsilon,\delta)$-JLM property.
@@
Consider the Block-Sparse Tensor Train sketch (4) with parameters $P,R$:
${\bm{\Omega}}_{\mathtt{BSTT}}=\frac{1}{\sqrt{P}}\begin{bmatrix}{\bm{\Omega}}_{\mathrm{GTT},1}\\ \vdots\\ {\bm{\Omega}}_{\mathrm{GTT},P}\end{bmatrix}$,
where ${\bm{\Omega}}_{\textrm{GTT},i}$ for $i\in[P]$ are iid realizations of a Gaussian TT sketch matrix (9) with uniform ranks $r_{0},\dots,r_{d-1}=R$. Then, the BSTT sketch matrix ${\bm{\Omega}}_{\mathtt{BSTT}}$ satisfies the Strong $(\varepsilon,\delta)$-JLM property provided
@@
Proof of Theorem 3.7 (OSE with BSTT Sketches).
@@
Proof of Corollary 3.9 (BSTT Sketch for RSVD).
From Proposition 6.12, Imported Lemma 6.4, Theorem 3.7, and Imported Lemma 6.10, the ${\bm{\Omega}}_{\mathrm{BSTT}}$ sketch allows to compute an $(1+\varepsilon)$-accurate low-rank approximation of rank $r$ provided:
@@
7 Oblivious Subspace Injection (OSI) property for Block-Sparse Tensor Train
In this section we present the proof of the OSI for Block-Sparse Tensor Train based on the Gaussian comparison technique introduced in [47] which estimates the minimum eigenvalue of random positive definite matrices.
@@ 7.1 Moments computation
allowing for a stochastic lower bound of $\lambda_{min}(\mathbf{Y})$ by $\lambda_{min}(\mathbf{Z})$ which can be studied by specialized methods for Gaussian matrices. To apply this program to the BSTT sketch, we need to estimate the 4th-order moments $\mathbb{E}\left[{\mathrm{Tr}\left({\mathbf{S}_{r}\mathbf{W}_{1}}\right)^{2}}\right]$ when ${\bm{\Omega}}_{1}$ is a Gaussian TT with uniform rank $R\geq 1$.
@@ 7.2 Gaussian Comparison Model
7.3 OSI property for the Block-Sparse Tensor Train Sketch
@@
$\sum_{\mathcal{I}\subsetneq[d]}\frac{\gamma_{\mathcal{I}}(R)}{\sqrt{n_{\mathcal{I}^{c}}}}\leq C_{\mathbf{Q}}(R)\leq\left(1+\frac{p_{\mathbb{F}}}{R}\right)^{d}-1\quad\implies\quad C_{\mathbf{Q}}(R)=\mathcal{O}\left(\frac{dp_{\mathbb{F}}}{PR}\right),$
@@
Proof of Theorem 3.10 (OSI with BSTT sketches).
Fix an arbitrary $\mathbf{Q}\in\mathbb{F}^{N\times r}$ with orthonormal columns. Using the same notation as in Section 7.2, the Gaussian comparison model to the Block-Sparse Tensor Train sketch writes:
@@
where $\mathbf{X}_{i}$ are iid copies of the Gaussian comparison model (20) for the the Gaussian TT sketch. From Theorem 2.3 in reference [47], for $\mathbf{Y}:=({\bm{\Omega}}\mathbf{Q})^{*}({\bm{\Omega}}\mathbf{Q})$ where $\bm{\Omega}$} is the BSTT sketch matrix,
@@
Further, by the sub-additivity of the weak variance:
@@
7.4 Randomized SVD guarantee with a Block-Sparse Tensor Train sketch
In this section, we establish probabilistic guarantees (Corollary 3.12) for a QB factorization (which may be further manipulated to construct a compact SVD) in the particular instances of the Block-Sparse Tensor Train sketch. As a first step, we show the following improvement on Lemma 2.1 in [8] with constant $C_{\delta}=\mathcal{O}(\alpha^{-1}\sqrt{d/(PR\delta})$ instead of $\mathcal{O}(1/\alpha\delta)$:
@@ 7.4 Randomized SVD guarantee with a Block-Sparse Tensor Train sketch
Fix two matrices with orthonormal columns $\mathbf{Q}\in\mathbb{F}^{N\times r}$ and $\mathbf{Q}_{\perp}\in\mathbb{F}^{N\times s}$ whose ranges are mutually orthogonal: $\mathbf{Q}^{*}\mathbf{Q}_{\perp}={\bm{0}}$, and pick $\delta>0$. Choose an arbitrary matrix $\mathbf{B}\in\mathbb{F}^{t\times s}$, and draw a random BSTT sketch matrix ${\bm{\Omega}}\in\mathbb{F}^{N\times PR}$ realizing an $(\alpha,\delta,r)$-OSI of the subspace spanned by the $r$ columns of $\mathbf{Q}$ as determined by (7).
@@
In turn, this allows to bound the error from the randomized SVD implemented with a Block-Sparse Tensor Train sketch.
Proof of Corollary 3.12 (BSTT Sketch with OSI for RSVD).
As with Theorem 2.2 in [8], the proof follows from the standard RSVD error bound: For a given matrix $\mathbf{A}\in\mathbb{F}^{k\times N}$ and BSTT sketch $\mathbf{A}{\bm{\Omega}}^{*}$, we compute the rank-$PR$ randomized factorization $\widehat{\mathbf{A}}:=\mathbf{Q}(\mathbf{Q}^{*}\mathbf{A})$ where $\mathbf{Q}\mathbf{Q}^{*}$ is the orthogonal projection onto the column space of $\mathbf{A}{\bm{\Omega}}^{*}$. Write the partitioned SVD
@@ 7.5 Analysis of Randomize-then-Orthogonalize
Let $\mathcal{A}\in\mathbb{F}^{n_{1}\times\dots\times n_{d}}$ be a tensor, $(r_{k})_{k\in[d-1]}\in\mathbb{N}^{d-1}$ be the target ranks, let $\mathbf{\Omega}_{\mathtt{BSTT}}=\mathcal{G}_{2}\bowtie\cdots\bowtie\mathcal{G}_{d}\in\mathbb{F}^{r_{1}\times n_{2}\times\dots\times n_{d}}$ be a Block-Sparse Tensor Train sketch with
@@ 8 Conclusions/future perspectives
In this work, we introduced the Block-Sparse Tensor Train sketch, a novel tensor sketching framework that naturally unifies and extends the Khatri-Rao and Gaussian TT approaches. We established theoretical embedding guarantees exhibiting linear scaling with respect to the tensor order, and as a direct consequence, we derived quasi-optimal error bounds for the randomized TT Rounding scheme and empirically demonstrated its computational efficiency across a range of relevant applications.
@@
Finally, we hope the BSTT sketch can be efficiently used for quantum chemistry applications, where it holds significant potential, since inherent physical symmetries induce block-sparse structures within the TT cores [5]. The primary theoretical challenge in this domain lies in determining how to optimally align the BSTT sketch with the natural block sparsity of these tensors.

### 13 Mar 2026 v2

@@ 6.2 Properties of Gaussian Tensor-Train Sketches
By Lemma 6.3, condition (13) ensures that the Gaussian matrices $\mathbf{G}_{j}$ satisfy the Strong $(\varepsilon/L\sqrt{d+1-j},\delta)$-JLM Property and by Imported Lemma 6.6, so do matrices $\mathbf{M}_{j}=\mathbf{I}_{n_{1}\cdots n_{j-1}}\otimes\mathbf{G}_{j}$. Finally, by Imported Lemma 6.7, the product ${\bm{\Omega}}=\mathbf{M}_{1}\cdots\mathbf{M}_{d}$ satisfies the Strong $(\varepsilon,\delta)$-JLM Property. ∎
@@
6.3 Properties of TTStack sketches
Now knowing that each row block of the TTStack sketch is indeed a subspace embedding for large enough $R$, we can proceed to showing that this unifying framework is also an OSE, where the number of blocks $P$ helps relax the constraints on the ranks $R$. To this effect, we show the following result:
@@ 6.3 Properties of TTStack sketches
Let ${\bm{\Omega}}_{1},\dots,{\bm{\Omega}}_{P}$ be iid realizations of a random matrix with the Strong $\bigl(\frac{\sqrt{P}\varepsilon}{4e^{2}},\delta\bigr)$-JLM property, then the random matrix ${\bm{\Omega}}=\frac{1}{\sqrt{P}}\begin{bmatrix}{\bm{\Omega}}_{1}\\ \vdots\\ {\bm{\Omega}}_{P}\end{bmatrix}$ has the Strong $(\varepsilon,\delta)$-JLM property.
@@
Consider the TTStack sketch (4) with parameters $P,R$:
${\bm{\Omega}}_{\mathtt{BSTT}}=\frac{1}{\sqrt{P}}\begin{bmatrix}{\bm{\Omega}}_{\mathtt{GTT},1}\\ \vdots\\ {\bm{\Omega}}_{\mathtt{GTT},P}\end{bmatrix}$,
where ${\bm{\Omega}}_{\mathtt{GTT},i}$ for $i\in[P]$ are iid realizations of a Gaussian TT sketch matrix (9) with uniform ranks $r_{0},\dots,r_{d-1}=R$. Then, the TTStack sketch matrix ${\bm{\Omega}}_{\mathtt{BSTT}}$ satisfies the Strong $(\varepsilon,\delta)$-JLM property provided
@@
Proof of Theorem 3.7 (OSE with TTStack Sketches).
@@
Proof of Corollary 3.9 (RSVD with TTStack.
From Proposition 6.12, Imported Lemma 6.4, Theorem 3.7, and Imported Lemma 6.10, the ${\bm{\Omega}}_{\mathrm{TTS}}$ sketch allows to compute an $(1+\varepsilon)$-accurate low-rank approximation of rank $r$ provided:
@@
7 Oblivious Subspace Injection (OSI) property for TTStack
In this section we present the proof of the OSI for TTStack based on the Gaussian comparison technique introduced in [47] which estimates the minimum eigenvalue of random positive definite matrices.
@@ 7.1 Moments computation
allowing for a stochastic lower bound of $\lambda_{min}(\mathbf{Y})$ by $\lambda_{min}(\mathbf{Z})$ which can be studied by specialized methods for Gaussian matrices. To apply this program to the TTStack sketch, we need to estimate the 4th-order moments $\mathbb{E}\left[{\mathrm{Tr}\left({\mathbf{S}_{r}\mathbf{W}_{1}}\right)^{2}}\right]$ when ${\bm{\Omega}}_{1}$ is a Gaussian TT with uniform rank $R\geq 1$.
@@ 7.2 Gaussian Comparison Model
7.3 OSI property for the TTStack Sketch
@@
$\sum_{\mathcal{I}\subsetneq[d]}\frac{\gamma_{\mathcal{I}}(R)}{\sqrt{n_{\mathcal{I}^{c}}}}\leq C_{\mathbf{Q}}(R)\leq\left(1+\frac{p_{\mathbb{F}}}{R}\right)^{d}-1\quad\implies\quad C_{\mathbf{Q}}(R)=\mathcal{O}\left(\frac{dp_{\mathbb{F}}}{R}\right),$
@@
Proof of Theorem 3.10 (OSI with TTStack sketches).
Fix an arbitrary $\mathbf{Q}\in\mathbb{F}^{N\times r}$ with orthonormal columns. Using the same notation as in Section 7.2, the Gaussian comparison model to the TTStack sketch writes:
@@
where $\mathbf{X}_{i}$ are iid copies of the Gaussian comparison model (20) for the the Gaussian TT sketch. From Theorem 2.3 in reference [47], for $\mathbf{Y}:=({\bm{\Omega}}\mathbf{Q})^{*}({\bm{\Omega}}\mathbf{Q})$ where $\bm{\Omega}$} is the TTStack sketch matrix,
@@
Further, by the sub-additivity of the weak variance and by Lemma 7.5
@@
7.4 Randomized SVD guarantee with a TTStack sketch
In this section, we establish probabilistic guarantees (Corollary 3.12) for a QB factorization (which may be further manipulated to construct a compact SVD) in the particular instances of the TTStack sketch. As a first step, we show the following improvement on Lemma 2.1 in [8] with constant $C_{\delta}=\mathcal{O}(\alpha^{-1}\sqrt{d/(PR\delta})$ instead of $\mathcal{O}(1/\alpha\delta)$:
@@ 7.4 Randomized SVD guarantee with a TTStack sketch
Fix two matrices with orthonormal columns $\mathbf{Q}\in\mathbb{F}^{N\times r}$ and $\mathbf{Q}_{\perp}\in\mathbb{F}^{N\times s}$ whose ranges are mutually orthogonal: $\mathbf{Q}^{*}\mathbf{Q}_{\perp}={\bm{0}}$, and pick $\delta>0$. Choose an arbitrary matrix $\mathbf{B}\in\mathbb{F}^{t\times s}$, and draw a random TTStack sketch matrix ${\bm{\Omega}}\in\mathbb{F}^{N\times PR}$ realizing an $(\alpha,\delta,r)$-OSI of the subspace spanned by the $r$ columns of $\mathbf{Q}$ as determined by (7).
@@
In turn, this allows to bound the error from the randomized SVD implemented with a TTStack sketch.
Proof of Corollary 3.12 (RSVD with TTStack OSI).
As with Theorem 2.2 in [8], the proof follows from the standard RSVD error bound: For a given matrix $\mathbf{A}\in\mathbb{F}^{k\times N}$ and TTStack sketch $\mathbf{A}{\bm{\Omega}}^{*}$, we compute the rank-$PR$ randomized factorization $\widehat{\mathbf{A}}:=\mathbf{Q}(\mathbf{Q}^{*}\mathbf{A})$ where $\mathbf{Q}\mathbf{Q}^{*}$ is the orthogonal projection onto the column space of $\mathbf{A}{\bm{\Omega}}^{*}$. Write the partitioned SVD
@@ 7.5 Analysis of Randomize-then-Orthogonalize
Let $\mathcal{A}\in\mathbb{F}^{n_{1}\times\dots\times n_{d}}$ be a tensor, $(r_{k})_{k\in[d-1]}\in\mathbb{N}^{d-1}$ be the target ranks, let $\mathbf{\Omega}_{\mathtt{BSTT}}=\mathcal{G}_{2}\bowtie\cdots\bowtie\mathcal{G}_{d}\in\mathbb{F}^{r_{1}\times n_{2}\times\dots\times n_{d}}$ be a TTStack sketch with
@@ 8 Conclusions/future perspectives
In this work, we introduced the TTStack sketch, a novel tensor sketching framework that naturally unifies and extends the Khatri-Rao and Gaussian TT approaches. We established theoretical embedding guarantees exhibiting linear scaling with respect to the tensor order, and as a direct consequence, we derived quasi-optimal error bounds for the randomized TT Rounding scheme and empirically demonstrated its computational efficiency across a range of relevant applications.
@@
Finally, we hope the TTStack sketch can be efficiently used for quantum chemistry applications, where it holds significant potential, since inherent physical symmetries induce block-sparse structures within the TT cores [5]. The primary theoretical challenge in this domain lies in determining how to optimally align the TTStack sketch with the natural block sparsity of these tensors.